OpenStudy (anonymous):

Please help: Find all solutions to the equation. sin^2(x) + sin (x) = 0

OpenStudy (anonymous):

ok u know that\[\sin 2x=2 \sin x\cos x\]so equation becomes\[2 \sin x\cos x+\sin x=0\]take out common factor\[\sin x(2\cos x+1)=0\]using zero product property\[\sin x=0\]or\[2\cos x+1=0\]

OpenStudy (anonymous):

but the equation is sin^2(x)

OpenStudy (anonymous):

oh sorry\[\sin^2 x+\sin x=0\]take out common factor\[\sin x(\sin x+1)=0\]easier :)

OpenStudy (anonymous):

\[\sin x=0\]or\[\sin x=-1\]

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

a little but i dont know how to get the solutions

OpenStudy (anonymous):

u start like this: Let x be sin(x) and x^2+x=0 x(x+1)=0 then, x=0; & x=-1 sin is at radian -3/4pi and pi

OpenStudy (anonymous):

u substitute x value in sin(x) and -1 is co-terminal (270) angle and 0